OpenStudy (anonymous):

Write the Riemann sum to find the area under the graph of the function f(x) = x2 from x = 1 to x = 5.

OpenStudy (anonymous):

@jim_thompson5910 just one minute let me work it out

OpenStudy (anonymous):

https://i.gyazo.com/794492d04bf957ce8ef52556a4f80504.png B?

jimthompson5910 (jim_thompson5910):

very good, it's just like the last one but now delta x = (b-a)/n = (5-1)/n = 4/n xi = a+i*(delta x) = 1+i*(4/n) = 1+(4i/n)

OpenStudy (anonymous):

Alright, my last question is how to do it in reverse. I'll post it in a minute.